Use the following results from a test for marijuana use, which is provided by a certain drug testing company, Among 141 subjects with positive test results, there are 30 false positive results. Among 157 negative results, there are 3 false negative results. Complete parts (a) through (Hint: Construct a table) 

a. How many subjects were included in the study? The total number of subjects in the study was _______ . 

b. How many subjects did not use marijuana? A lot of _______ subjects did not use marijuana 

c. What is the probability that a randomly selected subject did not use marijuana? The probability that a randomly selected subject did not use marijuana is _______  (Do not round the final answer. Then round to three decimal places as needed)
Answer #1

Marijuana use || Test result

Positive Negative Total
Yes 111 (True +ve) 3 (False -ve) 114
No 30 (False +ve) 154 (True -ve) 184
Total 141 157 298

From the given data, we constructed the table about.

a) There were total of 298 subjects included in the study.

b) Number of subjects who did not use marijuana = 30 + 154 = 184

c) Probability that a randomly selected subject did not use marijuana

= Number of subjects who did not use marijuana / Total number of subjects in study.

= 184/298 = 0.617